# Concurrent Fermionic Simulation Gate

### v1

@@ V Optimization and Results
$\begin{split}\tilde{G}^{{\rm eff}}_{110,020}=i&\alpha\Omega_{2}J_{0}\left(\frac{\beta\Omega_{1}}{\nu_{1}}\right)\left[J_{2}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)+J_{0}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)\right]\\ &\times\exp\left({i(\tilde{\Delta}_{110,020}-\nu_{2})t}\right)\end{split}$ (24)
@@
The effective Hamiltonian in this subspace which then can be written as:
$\begin{split}H=ig\hat{\Sigma}^{+}e^{i\Delta t}+H.c.\end{split}$ (27)
@@
$\begin{split}\Delta&=\Delta_{110,020}-\nu_{2}\\ g&=\alpha\Omega_{2}J_{0}\left(\frac{\beta\Omega_{1}}{\nu_{1}}\right)\left(J_{2}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)+J_{0}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)\right)\end{split}$ (28)
@@
$\begin{split}\tilde{H}&=U_{r}^{\dagger}HU_{r}-iU_{r}^{\dagger}\frac{\partial U_{r}}{\partial_{t}}\\ &=\frac{\Delta}{2}\hat{\Sigma}^{z}+g\hat{\Sigma}^{x}\end{split}$ (29)
@@
where $\Omega=\sqrt{(\frac{\Delta}{2})^{2}+g^{2}},n_{z}={\Delta}/{\sqrt{\Delta^{2}+4g^{2}}}$ and $n_{x}={2g}/{\sqrt{\Delta^{2}+4g^{2}}}$. Note that a part of conditional phase due to ZZ interaction is now absorbed in the rotating frame transformation $U_{r}$ [55].
@@
Eq.(32) is the main equation that we need for determining the optimal amplitude of the second drive. Note that once $\alpha,\beta$ and $\gamma$ are fixed in Eq.( 28) by a given undriven Hamiltionian, the sign of $g$ normally won’t change within a reasonable range of $\frac{\nu_{1,2}}{\Omega_{1,2}}$ unless we flip the signs of $\Omega_{1,2}$ or $\nu_{1,2}$. So we will only see a single branch in the numerical solution of $g$. The optimal amplitude shows up as a low-leakage trajectory around a high-leakage center in the 2D scan of $\nu_{2}$ and $\Omega_{2}$. Following the trajectory we have the lowest leakage ( complete cycle of $|110\rangle\rightarrow|020\rangle\rightarrow|110\rangle$) and also tunable cphase, see FIG.6.
Figure 6: Identify the optimal amplitude trajectory. The optimal trajectory is marked by the red dash line. (a) Analytical plot of P11 vs $\Omega_{2}$ and $\nu_{2}$ while the first drive is resonant with $|01\rangle\leftrightarrow|10\rangle$ transition: $\Omega_{1}=150$ MHz, $\nu_{1}=\Delta_{01,10}$. The optimal drive amplitude $\Omega_{2}$ is defined as such that P11 is maximum. (b) Numerical plot of P11 vs $\Omega_{2}$ and $\nu_{2}$. It is consistent with the analytical plot. (c) Numerical plot of leakage vs $\Omega_{2}$ and $\nu_{2}$. We calculate the average leakage out of the computational subspace when the initial states are $|010\rangle$, $|100\rangle$ and $|110\rangle$. The optimal drive amplitude $\Omega_{2}$ can also be defined as such to minimize the leakage.
@@
Figure 7: Optimal $\Omega_{2}$ vs $\nu_{2}$. 7(a) numerically determined optimal drive amplitude $\Omega_{2}$ by minimizing leakage and analytical optimal $\Omega_{2}$. The discrepancy on the edges is because $|\nu_{2}-\Delta_{110,020}|>\frac{2\pi}{t_{g}}$. No valid solution of effective coupling $g_{110,020}$ can be found to satisfy the condition that one and only one cycle of $|110\rangle\rightarrow|020\rangle\rightarrow|110\rangle$ is completed in the given gate time $t_{g}$. Therefore, only the $\nu_{2}$ and $\Omega_{2}$ in the middle area are relevant. 7(b) Analytically calculated optimal $\Omega_{2}$ vs $\nu_{2}$ with various $\Omega_{1}$.
VI Concurrency in the cfSim gate
Combining the results above, we can now write down the equations for the iswap angle $\theta$ and conditional phase $\varphi$ including the contribution from ZZ interaction and the constraint condition for $\Omega_{2}$ for a given set of $(t_{g},\Omega_{1},\nu_{2}$) with $\nu_{1}=\Delta_{100,010}$:
@@ VI Concurrency in the cfSim gate
$\begin{split}&\omega_{1}/2\pi=7.15\text{ GHz, }\omega_{2}/2\pi=7.6\text{ GHz,}\\ &\omega_{r}/2\pi=8.1\text{ GHz}\\ &\delta_{1}/2\pi=\delta_{2}/2\pi=-200\text{ MHz, }\delta_{r}=0\\ &g_{1r}/2\pi=g_{2r}/2\pi=120\text{ MHz, }g_{12}=0\\ &\phi_{1}^{d}=\phi_{2}^{d}=0\\ &t_{g}=100\text{ ns}\end{split}$ (36)
We first calculate the optimal drive amplitude $\Omega_{2}$ according to Eq.(32), see Fig. (13 a) in Appendix (D). We then use the analytically calculated $\Omega_{2}$ in the time-dependent Schrödinger equation and numerically solve the equation. Fig. (13 c) shows the leakage at the end of the gate. Within the valid area of Eq.(32), the leakage is suppressed below 1%. Note that in Eq.(35) we have assumed that the coupler $|R\rangle$ is in the ground state. This is valid because the coupler is far detuned from both qubits by more than 1 GHz. During the gate, no drive is resonant with any coupler excitation. So the coupler won’t be occupied during the gate execution. In Fig. (8), we plot the iswap angle $\theta$ and conditional phase $\varphi$ from numerical simulation and compare it to analytical results from Eq.(33). $\theta$ is tunable from 0 to $90^{\circ}$ and $\varphi$ is tunable from $-180^{\circ}$ to $180^{\circ}$. Both can be faithfully predicted by the analytical formulas. To quantify the effect of leakage, we calculate the gate fidelity using the formula[57]:
@@
The fidelity from the numerical simulation is shown in Fig. (9 a). In the fSim gate range, the fidelity is consistently above 99.5%. In most area it’s above 99.9%. The cross-talk effect can be clearly seen in the line cuts Fig. (9 c). The gate fidelity is lower when $\Omega_{1}$ is larger. This is most likely because of the beyond-RWA effect, which is left for future study.
@@
Figure 8: Continuous fSim gate. Both $\theta$ and $\varphi$ are continuously tunable. 8(a) Analytically calculated conditional phase vs $\Omega_{1}$ and $\nu_{2}$. The phase covers the full $2\pi$ range from $-\pi$ to $\pi$. The phase is dominantly determined by $\nu_{2}$. The vertical cut in the middle is because the range is chosen to be $[-\pi,\pi)$ in order to manifest the zero conditional phase trajectory. Along the vertical line $\varphi=\pi(-\pi)$. 8(b) Numerically calculated conditional phase vs $\Omega_{1}$ and $\nu_{2}$. It agrees with the analytical plot well. 8(c) Analytically calculated iswap angle $\theta$ vs $\Omega_{1}$ and $\nu_{2}$. $\theta$ is tunable from 0 to $90^{\circ}$. $\theta$ is mainly controlled by $\Omega_{1}$ with some modulation from $\nu_{2}$. A larger $\Omega_{1}$ is needed to achieve maximum $\theta$ when $\nu_{2}$ is resonant with $\Delta_{110,020}$. This is because when $\nu_{2}$ gets closer to $\Delta_{110,020}$, the corresponding optimal $\Omega_{2}$ becomes larger, which in turn makes $\frac{\Omega_{2}}{\nu_{2}}$ larger. The effective coupling $g_{010,100}$ is thus reduced, see Eq.(21) and Fig.(12). 8(d) Numerically calculated iswap angle $\theta$ vs $\Omega_{1}$ and $\nu_{2}$, consistent with analytical plot.
(a)Refer to caption(b)
Refer to caption(c)Refer to caption(d)
Figure 9: 9(a) Fidelity of the fSim gate. The dark area is out of the fSim gate range. Within the fSim gate range, the fidelity is around or above $9$(b) Fidelity along the $\nu_{2}$ axis while $\Omega_{1}$ is fixed. Three line cuts are shown. The fidelity domain are between 99.9% to 99.99% subject to qubit decoherence. 9(c) iswap angle $\theta$ vs $\Omega_{1}$ while $\nu_{2}$ is fixed. 9(d) Conditional phase $\varphi$ vs $\nu_{2}$ while $\Omega_{1}$ is fixed. The crosstalk effect on the conditional phase is small.
The gate fidelity can be further improved by including pulse-shaping. Here we choose to use flat-top Gaussian to demonstrate the effect of pulse-shaping. One crucial step in our numerical simulation is to find the optimal drive amplitude $\Omega_{2}$, see Eq.(33) and Eq.(34). This was calculated by using RWA when the drives are rectangular pulses. In the case of pulse-shaping, however, the RWA result is inaccurate because of the slow-changing rise-and-fall edges, especially when one wants to achieve high gate fidelity. In order to overcome this problem, we choose to make a two-level approximation and numerically optimize the drive amplitude in the effective two-level system, which is spanned by $|110\rangle$ and $|020\rangle$ in our case. One can also include more relevant levels accordingly. During the numerical optimization, only the pulse amplitude is varied. Other parameters, such as pulse width and rise-and-fall time, are fixed, which can also affect the gate fidelity. The pulse envelop to be optimized is therefore expressed as:
@@
$e(t)$ is the final pulse envelop and $f(t)$ is the standard flat top Gaussian envelop. $\Gamma$ is the parameter to be optimized. We find the optimal drive amplitude by requiring minimal leakage to $|020\rangle$, or equivalently maximal $|110\rangle$ occupation. The gate fidelity, iswap probability P10 and the conditional phase with pulse-shaping are shown in Fig. (14) in Appendix (E). We can see that the gate fidelity is improved by approximately one-order-of-magnitude. The iswap and cphase patterns are roughly unchanged. To illustrate the enhancement to fidelity, we show three specific fidelity trajectories versus $\nu_{2}$ in Fig. (10).
Nevertheless, we do find more discontinuities/jumps in the plots. This is because of the numerical optimization. During the optimization, especially when the drive frequency $\nu_{2}$ is approaching the boundaries of the valid area of fSim gates, the resulting optimization factors may not be continuous in contrast to the analytical formulas Eq.(33) and Eq.(34) used when there is no pulse-shaping. The discontinuities in the optimization factor can be seen clearly in Fig. (14) in Appendix (E).
Figure 10: Three specific line cuts of fidelities along the $\nu_{2}$ axis.
@@
VII $ZZ$-free iswap in the cfSim
@@
VIII Summary

### v2

@@ III Minimizing Leakage
$\begin{split}g_{2}=i&\alpha\Omega_{2}J_{0}\left(\frac{\beta\Omega_{1}}{\nu_{1}}\right)\left[J_{2}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)+J_{0}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)\right]\\ &\times\exp\left({i(\tilde{\Delta}_{110,020}-\nu_{2})t}\right)\end{split}$ (23)
@@
Note that Eq. (23) has similar structure as Eq. (21). The only difference is that because here we assume the drive frequency $\nu_{2}$ is slightly detuned from $\tilde{\Delta}_{110,020}$, there is an extra slow-varying phase factor $\exp\left({i(\tilde{\Delta}_{110,020}-\nu_{2})t}\right)$, details can be found in Appendix (C)
@@
The effective Hamiltonian in this subspace can then be written as:
$\begin{split}H=ig_{2}\hat{\Sigma}^{+}e^{i\Delta t}+H.c.\end{split}$ (26)
@@
$\begin{split}\Delta&=\Delta_{110,020}-\nu_{2}\\ g_{2}&=\alpha\Omega_{2}J_{0}\left(\frac{\beta\Omega_{1}}{\nu_{1}}\right)\left(J_{2}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)+J_{0}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)\right)\end{split}$ (27)
@@
$\begin{split}\tilde{H}&=U_{r}^{\dagger}HU_{r}-iU_{r}^{\dagger}\frac{\partial U_{r}}{\partial_{t}}\\ &=\frac{\Delta}{2}\hat{\Sigma}^{z}+g_{2}\hat{\Sigma}^{x}\end{split}$ (28)
@@
where $\Omega=\sqrt{(\frac{\Delta}{2})^{2}+g_{2}^{2}},n_{z}={\Delta}/{\sqrt{\Delta^{2}+4g_{2}^{2}}}$ and $n_{x}={2g_{2}}/{\sqrt{\Delta^{2}+4g_{2}^{2}}}$. Note that a part of conditional phase due to ZZ interaction is now absorbed in the rotating frame transformation $U_{r}$ [69].
@@
Eq. (31) is the main equation that we need for determining the optimal amplitude of the second drive. Note that once $\alpha,\beta$ and $\gamma$ are fixed in Eq. (27) by a given undriven Hamiltionian, the sign of $g_{1,2}$ normally won’t change within a reasonable range of $\frac{\nu_{1,2}}{\Omega_{1,2}}$ unless we flip the signs of $\Omega_{1,2}$ or $\nu_{1,2}$. So we will only see a single branch in the numerical solution of $g_{1,2}$. The optimal amplitude shows up as a low-leakage trajectory around a high-leakage center in the 2D scan of $\nu_{2}$ and $\Omega_{2}$. Following the trajectory we have the lowest leakage ( complete cycle of $|110\rangle\rightarrow|020\rangle\rightarrow|110\rangle$) and also tunable CPHASE, see Fig. 5.
Figure 5: Identify the optimal amplitude trajectory. The optimal trajectory is marked by the red dashed line. (a) Analytical plot of P11 vs $\Omega_{2}$ and $\nu_{2}$ while the first drive is resonant with $|01\rangle\leftrightarrow|10\rangle$ transition: $\Omega_{1}=150$ MHz, $\nu_{1}=\Delta_{01,10}$. The optimal drive amplitude $\Omega_{2}$ is defined as such that P11 is maximum. (b) Numerical plot of P11 vs $\Omega_{2}$ and $\nu_{2}$. It is consistent with the analytical plot. (c) Numerical plot of leakage vs $\Omega_{2}$ and $\nu_{2}$. We calculate the average leakage out of the computational subspace when the initial states are $|010\rangle$, $|100\rangle$ and $|110\rangle$. The optimal drive amplitude $\Omega_{2}$ can also be defined as such to minimize the leakage.
@@
Figure 6: Optimal $\Omega_{2}$ vs $\nu_{2}$. 6(a) numerically determined optimal drive amplitude $\Omega_{2}$ by minimizing leakage and analytical optimal $\Omega_{2}$. The discrepancy on the edges is because $|\nu_{2}-\Delta_{110,020}|>\frac{2\pi}{t_{g}}$. No valid solution of effective coupling $g_{110,020}$ can be found to satisfy the condition that one and only one cycle of $|110\rangle\rightarrow|020\rangle\rightarrow|110\rangle$ is completed in the given gate time $t_{g}$. Therefore, only the $\nu_{2}$ and $\Omega_{2}$ in the middle area are relevant. 6(b) Analytically calculated optimal $\Omega_{2}$ vs $\nu_{2}$ with various $\Omega_{1}$.
IV Concurrency in the cfSim gate
In this section, we will show how to achieve full range control of the iSWAP angle and conditional phase by using BPD in one run. The simultaneous drive scheme in principle also applies to other drive combinations. In Appendix (E), we show simulations of simultaneous resonant iSWAP and parametric CPHASE, where the two qubits are tuned into resonance and a parametric drive is also present for the 11-02/20 transition. In the main text, we focus on the BPD scheme. Combining the results above, we can now write down the equations for the iSWAP angle $\theta$ and conditional phase $\varphi$ including the contribution from ZZ interaction and the constraint condition for $\Omega_{2}$ for a given set of $(t_{g},\Omega_{1},\nu_{2}$) with $\nu_{1}=\Delta_{100,010}$:
@@ IV Concurrency in the cfSim gate
$\omega_{1}/2\pi$ $\omega_{2}/2\pi$ $\omega_{c}/2\pi$
7.15 GHz 7.6 GHz 8.5 GHz
$\delta_{1}/2\pi$ $\delta_{2}/2\pi$ $\delta_{c}/2\pi$
-200 MHz -200 MHz 0
$g_{1c}/2\pi$ $g_{2c}/2\pi$ $g_{12}/2\pi$
120 MHz 120 MHz 0
$\phi_{1}^{d}$ $\phi_{2}^{d}$ $t_{g}$
0 0 100 ns
We first calculate the optimal drive amplitude $\Omega_{2}$ according to Eq. (31). We then use the analytically calculated $\Omega_{2}$ in the time-dependent Schrödinger equation and numerically solve the equation. Within the valid area of Eq. (31), the leakage is suppressed below 1%. Note that in Eq. (34) we have assumed that the coupler $|R\rangle$ is in the ground state. This is valid because the coupler is far detuned from both qubits in GHz order. During the gate, no drive is resonant with any coupler excitation. So the coupler won’t be occupied during the gate execution. In Fig. (7), we plot the iSWAP angle $\theta$ and conditional phase $\varphi$ from numerical simulation and compare it to analytical results from Eq. (32). $\theta$ is tunable from 0 to $90^{\circ}$ and $\varphi$ is tunable from $-180^{\circ}$ to $180^{\circ}$. Both can be faithfully predicted by the analytical formulas. To quantify the effect of leakage, we calculate the gate fidelity using the formula[71]:
@@
The fidelity from the numerical simulation is shown in Fig. (8a). In the fSim gate range, the fidelity is consistently above 99.5%. In most area it’s above 99.9%. The cross-talk effect can be clearly seen in the line cuts Fig. (8c). The gate fidelity is lower when $\Omega_{1}$ is larger. This is most likely because of the beyond-RWA effect, which is left for future study. This fidelity is at the same level as that given by common two-qubit gate schemes such as iSWAP or CZ. Nevertheless, our flexible concurrent fSim gates can reduce circuit depth compared to the one-native-2Q-gate approach. This can improve the overall fidelity of a quantum circuit.
Figure 7: Continuous fSim gate. Both $\theta$ and $\varphi$ are continuously tunable. 7(a) Analytically calculated conditional phase vs $\Omega_{1}$ and $\nu_{2}$. The phase covers the full $2\pi$ range from $-\pi$ to $\pi$. The phase is dominantly determined by $\nu_{2}$. The vertical cut in the middle is because the range is chosen to be $[-\pi,\pi)$ in order to manifest the zero conditional phase trajectory. Along the vertical line $\varphi=\pi(-\pi)$. 7(b) Numerically calculated conditional phase vs $\Omega_{1}$ and $\nu_{2}$. It agrees with the analytical plot well. 7(c) Analytically calculated iSWAP angle $\theta$ vs $\Omega_{1}$ and $\nu_{2}$. $\theta$ is tunable from 0 to $90^{\circ}$. $\theta$ is mainly controlled by $\Omega_{1}$ with some modulation from $\nu_{2}$. A larger $\Omega_{1}$ is needed to achieve maximum $\theta$ when $\nu_{2}$ is resonant with $\Delta_{110,020}$. This is because when $\nu_{2}$ gets closer to $\Delta_{110,020}$, the corresponding optimal $\Omega_{2}$ becomes larger, which in turn makes $\frac{\Omega_{2}}{\nu_{2}}$ larger. The effective coupling $g_{010,100}$ is thus reduced, see Eq. (20) and Fig. (4). 7(d) Numerically calculated iSWAP angle $\theta$ vs $\Omega_{1}$ and $\nu_{2}$, consistent with analytical plot.
@@
Figure 8: 8(a) Fidelity of the fSim gate. The dark area is out of the fSim gate range. Within the fSim gate range, the fidelity is around or above 99.9%. 8(b) Fidelity along the $\nu_{2}$ axis while $\Omega_{1}$ is fixed. Three line cuts are shown. The fidelity domain is between 99.9% to 99.99% , subject to qubit decoherence. 8(c) iSWAP angle $\theta$ vs $\Omega_{1}$ while $\nu_{2}$ is fixed. 8(d) Conditional phase $\varphi$ vs $\nu_{2}$ while $\Omega_{1}$ is fixed. The crosstalk effect on the conditional phase is small.
The gate fidelity can be further improved by including pulse-shaping. Here we choose to use flat-top Gaussian to demonstrate the effect of pulse-shaping. One crucial step in our numerical simulation is to find the optimal drive amplitude $\Omega_{2}$, see Eq. (32) and Eq. (33). This was calculated by using RWA when the drives are rectangular pulses. In the case of pulse-shaping, however, the RWA result is inaccurate because of the slow-changing rise-and-fall edges, especially when one wants to achieve high gate fidelity. In order to overcome this problem, we choose to make a two-level approximation and numerically optimize the drive amplitude in the effective two-level system, which is spanned by $|110\rangle$ and $|020\rangle$ in our case. One can also include more relevant levels accordingly. During the numerical optimization, only the pulse amplitude is varied. Other parameters, such as pulse width and rise-and-fall time, are fixed, which can also affect the gate fidelity. The pulse envelope to be optimized is therefore expressed as:
@@
$e(t)$ is the final pulse envelope and $f(t)$ is the standard flat top Gaussian envelope. $\Gamma$ is the parameter to be optimized. The flat top Gaussian $f(t)$ is defined as:
$f(t)=\begin{cases}\exp(-\frac{(t-\tau_{1})^{2}}{2\sigma^{2}})&\text{, }0<t<\tau_{1}\\ 1&\text{, }\tau_{1}<t<\tau_{1}+\tau_{2}\\ \exp(-\frac{(t-\tau_{1}-\tau_{2})^{2}}{2\sigma^{2}})&\text{, }\tau_{1}+\tau_{2}<t<2\tau_{1}+\tau_{2}\end{cases}$ (37)
We find the optimal drive amplitude by requiring minimal leakage to $|020\rangle$, or equivalently maximal $|110\rangle$ occupation. We find that the gate fidelity is improved by approximately one-order-of-magnitude. The iSWAP and CPHASE patterns are roughly unchanged. To illustrate the enhancement to fidelity, we show three specific fidelity trajectories versus $\nu_{2}$ in Fig. (9).
Nevertheless, we do find more discontinuities/jumps in the plots. This is because of the numerical optimization. During the optimization, especially when the drive frequency $\nu_{2}$ is approaching the boundaries of the valid area of fSim gates, the resulting optimization factors may not be continuous in contrast to the analytical formulas Eq. (32) and Eq. (33) used when there is no pulse-shaping.
Figure 9: Three specific line cuts along the $\nu_{2}$ axis of fSim gate fidelities with flat-top Gaussian pulse-shaping. The fidelities are improved by almost one order of magnitude. We use larger drive amplitudes $\Omega_{1}$ than in Fig. (8b) because it requires stronger drive to achieve the same iSWAP angles after adding pulse-shaping.
@@
V $ZZ$-free iSWAP in the cfSim
@@
VI Discussion and outlook
We have shown that our BPD scheme can efficiently and independently control iSWAP operation and CPHASE operation with high fidelity. Because we focus on the unitary dynamics behind BPD drive, we have left out the crucial topic of optimal gate time considering trade-off between unitary errors and decoherence.
Unitary errors such as leakage usually can be suppressed by prolonging gate time, but the gate time upper bound is set by T1 and T2 because at some point decoherence becomes the dominant error source. Faster gates can reduce decoherence while they suffer from increased unitary errors. In our analysis, we have mainly focused on reducing leakage and neglected most of non-RWA effects. Our numerical simulation of Hamiltonian Eq. (14) is exact. There is no extra approximation used other than discretizing the time-dependent Hamiltonian into piecewise constant Hamiltonians with small time steps and truncating the Hilbert space to 100 levels ( 5 levels for each transmon and 4 levels for the coupler). In this way, we partially capture non-RWA and other nonlinear effects. However, as our starting point is the RWA Hamiltonian Eq. (14), most of the non-linear effects are already approximated away from the very beginning. To fully address all strong drive effects, one needs to use the full circuit Hamiltonian as in Eq. (2) and Eq. (3). In fact, in another comprehensive study of strong drives on transmons, it is found that strong driving can ionize transmons and the system enters a regime of chaos[72]. Here we provide some preliminary simulation and study of the non-RWA and other non-linear effects without using the full circuit Hamiltonian model.
Instead of the RWA Hamiltonian Eq. (14), we include counter-rotating terms to get a non-RWA Hamiltonian:
$\begin{split}&H=H_{0}+H_{d},\\ &H_{0}=\sum_{\alpha=1,2,c}\omega_{\alpha}\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}+\frac{\delta_{j}}{2}\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}(\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}-1)\\ &\quad\quad\quad+\sum_{j,k=\{1,2\}}g_{jc}(\hat{a}^{\dagger}_{j}+\hat{a}_{j})(\hat{a}^{\dagger}_{c}+\hat{a}_{c}),\\ &H_{d}=\sum_{j=1,2}\Omega_{j}\sin(\nu_{j}t+\phi^{d}_{j})\hat{a}^{\dagger}_{j}\hat{a}_{j}.\end{split}$ (40)
We perform the same cfSim simulation as in Fig. (7) and (8). We find that in the middle area where the drives are strong, gate fidelity drops to around 99%, see Fig. (11). This is due to the extra non-RWA terms. We have used the same amplitude formula in Eq. (31), which is analytically derived from the RWA Hamiltonian. In the middle area, where $\nu_{2}$ is close to $\Delta_{110,020}$, the required drive amplitude $\Omega_{2}$ is large. This gives a stronger non-RWA effect, resulting in the fidelity drop. Luckily, much of this non-RWA error can be corrected via pulse-shaping. We again use the same pulse-shaping and optimization procedure as in Eq. (36) and perform the same simulation as in Fig. (9). The fidelity results are shown in Fig. (12). The fidelity loss is recovered after introducing a smooth envelope and using numerically optimized amplitudes. The non-RWA effects also change the iSWAP angle pattern while the conditional phases remain mostly unchanged. In general, iSWAP rate is reduced due to the non-RWA terms. Simulation results are included in Appendix (D).
(a)Refer to caption(b)
[…]
(a)Refer to caption(b)
[…]
Apart from non-RWA, another major nonlinear effect is the higher harmonics from drives[55]. We have assumed the drive Hamiltonian takes the simple form of $H_{d}=\sum_{j=1,2}\Omega_{j}\sin(\nu_{j}t+\phi^{d}_{j})\hat{a}^{\dagger}_{j}\hat{a}_{j}$. A harmonic drive in the full circuit Hamiltonian Eq. (3) would generate higher-order tones, especially when the drive is strong. In theory, this nonlinearity can be reversed by pulse-shaping, too. Suppose there is a relation between the drive in the Fock space (Eq. (14)) and the drive in the full circuit Hamiltonian Eq. (3), one can reverse engineer the drive envelope so that the drive in Eq. (14) is single-tone. This, of course, can be challenging to implement in experiments because the relation between the two is unknown in general. As an example, if the relation between two drives is given by $\omega(t)=\omega_{0}\sqrt{\varphi(t)/2}$[54, 73, 55], one can substitute in $\omega(t)=\Omega\sin(\nu t)$ to get $\varphi(t)=2\left(\frac{\Omega}{\omega_{0}}\right)^{2}\sin^{2}(\nu t)$. This would then give the desired single-tone drive in Eq. (14).
Dealing with the strong drive effects is crucial for parametric drives. This is because, compared to gates without using parametric modulations, parametric gates are usually slower under the same hardware conditions. One solution is to find novel schemes to increase parametric gate speed without entering the strong drive regime. For instance, there has been a promising attempt at using extra levels to speed up parametric transitions from Chalmers University of Technology[74, 75].
For future work, to get an accurate comparison with experiments and fully assess gate fidelities, full circuit Hamiltonians and open system simulations will be needed. This would help to understand what is the optimal gate time under trade-off between unitary errors and decoherence. Another direction is to apply the same simultaneous-driving principles to other gate setups. One of the easiest adaptations is to apply a parametric drive on top of a pair of resonant qubits so that iSWAP transition is activated by the resonance condition and CPHASE operation is controlled via the parametric drive. We provide some preliminary simulations in Appendix. E.
VII Summary